# Specialized Foundation Models Struggle to Beat Supervised Baselines

### v1

@@ 1 Introduction
Driven by this success, the foundation model approach has been adapted to various specialized domains, which we define to be ML application areas—e.g. genomics, satellite imaging, and time series—whose data modalities lie outside those of classical AI tasks, i.e. natural images and text. These domains have seen the introduction of many new FMs claiming to leverage large, domain-specific pretraining datasets to achieve breakthrough performance on downstream tasks (Dalla-Torre et al., 2023; Nguyen et al., 2024; Zhou et al., 2023b; Avsec et al., 2021; Ji et al., 2021; Fuller et al., 2023; Cong et al., 2022; Mendieta et al., 2023). These claims underlie our study’s motivating question:
@@
Figure 1: Across three domains—genomics, satellite imaging, and time series—specialized FMs fail to significantly improve upon tuned supervised learning despite using two-to-five orders of magnitude more data. In contrast, breakthrough FMs such as BERT dramatically outperformed supervised baselines in NLP (top left), causing the field to switch to fine-tuning as the default approach. For each domain we plot total pretraining and fine-tuning data used vs. the mean improvement across tasks over the supervised state-of-the-art. Specifics of our evaluations on the last three domains are in Section 4, while the NLP results are derived from the GLUE benchmark (Wang et al., 2019). Note that in the x-axis of the top left figure we ignore tokens used to pretrain word embeddings.
@@
We answer our motivating question by considering a reasonably representative set of three specialized domains—chosen according to the presence of multiple FMs and a standard set of evaluation tasks—and comparing their performance on those tasks with that of a traditional supervised learning workflow. As depicted in Figure 2, the latter is a model development, hyperparameter tuning, and training process in which all steps use only data from the target task, in contrast to the FM workflow, which uses vast amounts of pretraining data. By leveraging model selection tools ranging from classical information criteria to cutting-edge architecture search, we build automated pipelines that efficiently develop and train strong supervised models on over fifty tasks across three distinct domains.
Our main result is negative: we find that, despite being pretrained on massive datasets, specialized FMs struggle and very often fail to outperform models trained exclusively on downstream task data with traditional supervised learning (c.f. Figure 1). Specifically, we show that lightly adapted convolutional neural network (CNN) architectures such as wide ResNet and UNet attain state-of-the-art on the Nucleotide Transformer benchmark in genomics and match the latest pretrained satellite FMs on downstream classification. Furthermore, we show that tuned linear auto-regression (AR) matches or outperforms every open-source time series FM on a standard suite of seven forecasting tasks, despite using four or more orders of magnitude fewer parameters and data.
These results demonstrate that genomics, satellite imaging, and time series have not yet had their “BERT moment” (Devlin et al., 2019), i.e. these domains have not yet pretrained FMs that dominate traditional supervised approaches. This is despite the fact that all them have BERT-scale [^1] Models with 100M+ parameters trained on 100x or more data than supervised tasks in the domain are given. FMs and the fact that many of them are already witnessing a shift towards not comparing with supervised approaches, as was seen in natural language processing (NLP) post-BERT. More broadly, since these domains are among the most high-profile areas with specialized FMs, our results challenge the prevailing assumption that pretrained models yield superior performance. They also reinforce the need for robust and well-tuned baselines, with surprising findings such as (a) simply tuning kernel sizes and dilation rates in standard CNN backbones dominates a genomics classification benchmark and (b) rescuing the century-old AR forecaster from obsolescence is as easy as considering lookback parameters larger than five and training on a GPU. To facilitate ongoing research in these and other domains, we make code associated with both our CNN-tuning pipeline (DASHA [^2] https://github.com/ritvikgupta199/DASHA) and our AR-on-GPU workflow (Auto-AR [^3] https://github.com/Zongzhe-Xu/AutoAR) publicly available.
@@ 2 Related work
2.1 Specialized foundation models
Collectively our three target domains have more than twenty-five FMs, many developed via the “lift-and-shift” approach—borrowing terminology from Rolf et al. (2024)—in which techniques from core AI areas such as vision and language processing are applied with modest tailoring to specialized domains. In particular, many methods are built on out-of-domain models such as BERT, Swin, and Hyena (Ji et al., 2021; Mendieta et al., 2023; Nguyen et al., 2024; Shen et al., 2024a), with adaptations such as specialized tokenizations, embeddings, and model modifications for handling domain-specific considerations such long-range dependencies (Dalla-Torre et al., 2023; Zhou et al., 2023b; Das et al., 2023; Cohen et al., 2024; Shen et al., 2024b) or multispectral data (Cong et al., 2022).
@@ 2.1 Specialized foundation models
While the “lift-and-shift” approach can often be useful or at least a good starting point, its widespread use underlines the need for strong in-domain baselines to make sure that the combination of out-of-domain tooling and massive pretraining data is actually helpful. Such comparisons are not always conducted, e.g. the satellite FM SatMAE (Cong et al., 2022) is compared to ImageNet-initialized and randomly initialized ResNet-50 (He et al., 2015), while most of the time series FMs we consider only do a full comparison to one linear baseline, DLinear (Zeng et al., 2023). While this can sometimes be justified—e.g. in the case of NLP post-BERT—our results suggest that for now, specialized FMs should still compare to in-domain supervised model development.
@@
Lastly, we note that we are not the first to take a somewhat critical look at specialized FMs. For example, Yang et al. (2024) questioned the dominance of Transformers for protein sequence FMs by showing that convolutions could do just as well, which is related to our discovery that (supervised) CNNs were competitive with (largely Transformer-based) genomics FMs. Another study by Kedzierska et al. (2023) found that training an in-domain generative model could outperform pretraining one in single-cell biology applications, a finding generalized by our own results, which demonstrate that the underperformance of specialized FMs relative to in-domain training may be a broader trend across multiple domains. In the time series domain, Tan et al. (2024) show that the popular approach of fine-tuning text-pretrained LLMs on time series tasks often underperforms supervised models (in their case randomly initialized attention); our results generalize this to other time series FMs and use an even simpler supervised model (AR) to do so.
2.2 Specialized baselines
@@
Both of the automated supervised learning pipelines we develop are heavily influenced by successful in-domain model development. In particular, the NAS-based pipeline we use to achieve our results in genomics and satellite imaging is inspired by the success of the human-driven specification of kernel sizes and dilation rates in successful architectures like TCN (Lea et al., 2016) and ConvNeXt (Liu et al., 2022). At the same time, for time series our approach is based upon a well-tuned GPU implementation of perhaps the most basic forecasting model, AR.
2.3 AutoML for specialized domains
@@
While often evaluated on domains such as vision, automated techniques have long been used in specialized domains as well. An important example is Auto-ARIMA (Hyndman & Khandakar, 2008) for time series, although it has been found to underperform on the specific suite of tasks we consider (Challu et al., 2022). However, to avoid requiring significant expertise in any one domain, we also make use of AutoML methods developed specifically for diverse tasks (Roberts et al., 2021b; Shen et al., 2023), in particular the NAS method DASH (Shen et al., 2022) that can discover good kernel sizes and dilation rates for a CNN backbone faster than it can be trained from scratch.
@@ 3 Methodology
While it is standard to automate the last two steps of the procedure, model development is typically done by hand and so is difficult to do for fifty tasks across three domains. As a result, we settle for approximating the traditional supervised learning workflow by simulating the model development component using neural architecture search. To ensure fair comparison and reduce computational costs, we restrict ourselves to low-fidelity NAS methods that return an architecture in less time than it takes to train it. The results we obtain using NAS can therefore be viewed as lower bounds on the performance of supervised learning, as the model development might be significantly improved using less-heuristic or human-driven architecture design.
In the remainder of this section we detail how we handle the different steps of the supervised learning pipeline. Note that our NAS-dependent supervised workflow (DASHA)—which we cover in the first part of this section—yields our main results for genomics and satellite imaging but not for time series; in that domain we find its performance to be less competitive. There we instead focus on an even simpler approach based on linear auto-regression, whose model development and tuning we describe in the second subsection.
@@ 3.1 DASHA: Simulating the supervised workflow using NAS
As described in Algorithm 1, we augment the existing DASH approach in two ways: (1) trying more than one CNN backbone (e.g. both wide ResNet and UNet (Ronneberger et al., 2015)) and (2) using the well-known hyperparameter tuner ASHA (Li et al., 2020) to configure architecture-specific training settings. This combination gives our workflow its name. Following the NAS and hyperparameter tuning stages, we train the discovered architecture with the selected configuration on the target data. Further details, including the resources given to the three steps of the pipeline and the exact search spaces used by DASH and ASHA, are provided in Appendix B.1. Note that, while our focus is on data-efficient baselines, we do ensure that the entire workflow is never substantially more computationally expensive than fine-tuning an FM.
@@ 4.1 Genomics
We begin our investigation in the genomics domain, which has witnessed the development of numerous FMs, including the early Enformer Avsec et al. (2021), the DNABERT series (Ji et al., 2021; Zhou et al., 2023b), the HyenaDNA family (Nguyen et al., 2024), GENA-LM [^4] We compare to GENA-LM in Appendix C, as its reported metrics differ from the NT benchmark. (Fishman et al., 2024), the recent Caduceus family (Schiff et al., 2024), and the NT family (Dalla-Torre et al., 2023); The latter includes models with up to 2.5B parameters. To evaluate them, we consider the Nucleotide Transformer (NT) benchmark of Dalla-Torre et al. (2023), which contains eighteen tasks in three main categories: regulatory elements, RNA production, and histone modification. We use this benchmark because of its diversity and because it has been evaluated on by all of the aforementioned FMs, allowing us to include eight of them in the comparison.
Our numbers for these models are taken from Dalla-Torre et al. (2023, Supplementary Table 6); Following Dalla-Torre et al. (2023, Supplementary Table 5), We use F1 score and accuracy to evaluate a subset of regulatory elements and RNA production tasks, and we use Matthew’s Correlation Coefficient (MCC) as the main metric for evaluation on the remaining datasets.
4.1.1 Baselines
CNNs have long been used for genomics tasks (Avsec et al., 2020; Zhou & Troyanskaya, 2015) and so constitute natural supervised baselines; in particular we include 1D variants of Wide ResNet (WRN) and UNet, which we find perform better than some domain-specific CNNs. We use these same two backbones as the candidate CNNs tuned and selected from by our DASHA workflow.
Model Model Pretraining Avg. Avg. Mean Median
@@ 4.1.1 Baselines
NT-Multispecies (2.5B) 2.5B 174B 0.697 4.08 40.51 45.52
DNABERT-2 117M 32.5B 0.680 6.88 38.65 43.59
@@
Supervised Methods
Wide ResNet 2.0M 0 0.694 6.83 37.16 43.08
UNet 4.5M 0 0.68 7.78 38.67 42.69
@@
Table 1: Aggregate performance on genomics tasks, showing that our supervised workflow (DASHA) attains state-of-the-art on the NT benchmark, outperforming all FMs according to most measures while using no pretraining data and oftentimes many fewer parameters. For Mean/Median %Imp., we report percentage improvement over the Raw Probe baseline from Dalla-Torre et al. (2023), and for DASHA the model size refers to the largest configuration across tasks. “-” indicates unknown quantities.
4.1.2 Results
Our genomics results are displayed in Table 1, which shows that our supervised workflow (DASHA) consistently outperforms all FMs across all aggregate metrics. As discussed in Appendix C, our strong performance is driven in large part by outstanding performance on the histone modification tasks (c.f. Table 9). The more detailed results also highlight the importance of considering diverse baselines, with Wide ResNet usually being the selected architecture but UNet performing significantly better for promoter and splice site classification tasks. Overall, DASHA arguably sets a new state-of-the-art on the NT benchmark and certainly demonstrates that supervised methods remain quite competitive in genomics, despite the availability of massive pretraining datasets.
@@ 4.2 Satellite imaging
While they do not get as large as those in genomics, numerous BERT-scale FMs have also been introduced for satellite imaging, including SeCo (Manas et al., 2021), the SatMAE family (Cong et al., 2022), the CROMA family (Fuller et al., 2023), GFM (Mendieta et al., 2023), Scale-MAE (Reed et al., 2023), Satlas (Bastani et al., 2023), Prithvi (Jakubik et al., 2023), and SkySense (Guo et al., 2024). Because our evaluation includes GeoBench (Lacoste et al., 2024), a recently introduced satellite benchmark that has not been considered by many of these FMs, we obtain all results using our own fine-tuning; therefore we only consider a restricted subset of top-performing, open-source, and compatibly-formatted models. In all cases we use the fine-tuning workflow suggested by the authors of each FM plus some automated hyperparameter tuning; note that even with the original code and extra tuning our reproductions on previous benchmarks systematically underperformed results reported in the original works. We take our tasks mainly from GeoBench’s five classification tasks and then add four additional tasks—BigEarthNet (Sumbul et al., 2019), EuroSAT (Helber et al., 2019), Canadian Cropland (Jacques et al., 2023), and fMoW-Sentinel (Cong et al., 2022)—that are commonly used to evaluate other FMs. [^5] In Appendix C we report results when excluding tasks where missing channels may affect performance. As we focus on classification—sometimes with multiple labels—we report top-1 accuracy or mAP as appropriate.
@@
Foundation Models
SatMAE-Base 85.6M 700K 76.99 6.22 5.27 3.59
SatMAE-Large 303M 700K 77.75 4.5 6.52 4.62
GFM 86.8M 1.3M 77.18 5.56 5.77 4.08
SwinT-Base 86.8M 14M 76.69 5.28 4.86 1.43
CROMA-Base 90.6M 2M 77.39 4.33 5.85 4.22
CROMA-Large 312M 2M 78.03 3.33 6.90 6.09
Supervised Methods
ResNet50 23.5M 0 73.76 8.34 0.30 00.07
Wide ResNet 17.2M 0 73.97 8.22 0.00 0.00
UNet 17.3M 0 75.73 5.89 3.01 1.07
DASHA (our workflow) 32.4M 0 77.85 3.33 6.67 5.16
Table 2: Aggregate performance on satellite imaging tasks, demonstrating that a supervised learning workflow (DASHA) can match the performance of state-of-the-art specialized FMs, all while using no pretraining data and having two-to-ten times fewer parameters. For Mean/Median %Imp. we report percentage improvement over a vanilla Wide ResNet, and for DASHA the model size refers to the largest configuration across tasks.
4.2.1 Baselines
@@
4.2.2 Results
Table 2 shows that our supervised workflow attains the best or second-best performance across all aggregate metrics and is only ever slightly outperformed by CROMA-large. Notably, unlike in genomics, the FMs here consistently outperform CNN backbones, likely because the associated papers compare to them as baselines. However, the frequently superior performance of DASHA suggests that domain-aware model development would yield good supervised models in this field. Another contrast with genomics is that the larger versions of the FMs consistently attain superior performance here, suggesting they are making at least somewhat effective use of the pretraining data. Nevertheless, that this improvement can also be attained by DASHA, which uses no pretraining and produces a model that is ten times smaller, suggests that there remains significant room for improvement.

### v2

@@ 1 Introduction
Driven by this success, the foundation model approach has been adapted to various specialized domains, which we define to be ML application areas—e.g. genomics, satellite imaging, and time series—whose data modalities lie outside those of classical AI tasks, i.e. natural images and text. Such domains have seen the introduction of many new FMs claiming to leverage large pretraining datasets to achieve breakthrough performance on downstream tasks (Dalla-Torre et al., 2023; Nguyen et al., 2024; Zhou et al., 2023b; Avsec et al., 2021; Ji et al., 2021; Fuller et al., 2023; Cong et al., 2022; Mendieta et al., 2023). These claims underlie our study’s motivating question:
@@
Figure 1: Across genomics, satellite imaging, and time series, specialized FMs fail to significantly improve upon supervised learning despite using two-to-five orders of magnitude more data. In contrast, breakthrough FMs such as BERT dramatically outperformed supervised NLP baselines (top left), causing the field to switch to fine-tuning as the default approach. For each domain we plot total pretraining and fine-tuning data vs. mean improvement over the supervised state-of-the-art. The NLP results are from the GLUE benchmark (Wang et al., 2019) while evaluations of the last three domains are in Section 4. Note that for the NLP x-axis we ignore word embedding pretraining tokens.
@@
We answer our motivating question by considering a representative set of three specialized domains—chosen according to the presence of multiple FMs and a standard set of evaluation tasks—and comparing their performance on those tasks with that of a traditional supervised learning workflow. The latter is a three step process (i.e., model development, hyperparameter tuning, training) in which all steps use only data from the target task, in contrast to the FM workflow, which uses vast amounts of pretraining data (c.f. Figure 2). By leveraging model selection tools ranging from classical information criteria to cutting-edge architecture search, we build automated pipelines that efficiently develop and train strong supervised models on over fifty tasks across three distinct domains.
Our main result is negative: despite pretraining on massive data, specialized FMs struggle and often fail to outperform models trained solely on downstream task data with traditional supervised learning (c.f. Figure 1). Specifically, lightly adapted convolutional neural network (CNN) architectures such as wide ResNet and UNet attain state-of-the-art on the Nucleotide Transformer benchmark in genomics and match the latest satellite FMs on downstream classification. Furthermore, tuned linear auto-regression (AR) matches or outperforms every open-source time series FM on a standard suite of forecasting tasks, despite using four or more orders of magnitude fewer parameters and data.
These findings demonstrate that genomics, satellite imaging, and time series have not yet had their “BERT moment” (Devlin et al., 2019), i.e. these domains have not yet pretrained FMs that dominate traditional supervised approaches. This is despite the fact that all of them have BERT-scale [^1] Models with 100M+ parameters trained on 100x or more data than supervised tasks in the domain are given. FMs and many are already witnessing a shift towards not comparing with supervised approaches, as was seen in natural language processing (NLP) post-BERT. More broadly, since these domains are among the most high-profile areas with specialized FMs, our results suggest that practitioners should still consider supervised models in addition to FMs. They also reinforce the need for robust and well-tuned baselines, with surprising findings such as (a) simply tuning kernel sizes and dilation rates in standard CNNs dominates a genomics benchmark and (b) rescuing the century-old AR forecaster from obsolescence is as easy as considering lookback parameters larger than five and training on a GPU. To facilitate ongoing research in these and other domains, we make code associated with both our CNN-tuning pipeline (DASHA) and our AR-on-GPU workflow (Auto-AR) publicly available. [^2] Available at https://github.com/ritvikgupta199/DASHA and https://github.com/Zongzhe-Xu/AutoAR.
@@ 2 Related work
Specialized foundation models:
@@
Collectively our three target domains have more than thirty FMs, many developed via the “lift-and-shift” approach—borrowing terminology from Rolf et al. (2024)—in which techniques from core AI areas such as vision and language processing are applied with modest tailoring to specialized domains. In particular, many methods are built on out-of-domain models such as BERT, Swin, and Hyena (Ji et al., 2021; Mendieta et al., 2023; Nguyen et al., 2024; Shen et al., 2024b), with adaptations including specialized tokenizations, embeddings, and model modifications for handling domain-specific considerations like long-range dependencies (Dalla-Torre et al., 2023; Zhou et al., 2023b; Shen & Yang, 2021; Das et al., 2023; Shen et al., 2024a; c; 2025; Cohen et al., 2024; Liang et al., 2025) or multispectral data (Cong et al., 2022). While the “lift-and-shift” approach can often be useful or at least a good starting point, its widespread use underlines the need for strong in-domain baselines to make sure that the combination of out-of-domain tooling and massive pretraining data is actually helpful. Such comparisons are not always conducted, e.g. the satellite FM SatMAE (Cong et al., 2022) is compared to ImageNet-initialized and randomly initialized ResNet-50 (He et al., 2015), while most of the time series FMs we consider only do a full comparison to one linear baseline, DLinear (Zeng et al., 2023). This can sometimes be justified—e.g. in the case of NLP post-BERT—but our results suggest that for now, specialized FMs should still compare to in-domain supervised model development.
@@
We note that we are not the first to take a critical look at specialized FMs. For example, Yang et al. (2024) questioned the dominance of Transformers for protein sequence FMs by showing that convolutions could do just as well, which is related to our discovery that (supervised) CNNs were competitive with (largely Transformer-based) genomics FMs. Another study by Kedzierska et al. (2023) found that training an in-domain generative model could outperform pretraining in single-cell biology applications. In satellite imaging, Jakubik et al. (2023) showed that a supervised UNet performs competitively with an FM on a single non-standard task, while Xiong et al. (2024) found that supervised models sometimes outperform linearly probed FMs. Corley et al. (2024) further demonstrated that ImageNet-pretraining can be effective in this domain when properly tailored. Finally, in the time series domain, Tan et al. (2024) show that the popular approach of fine-tuning text-pretrained LLMs on time series tasks often underperforms supervised models (in their case randomly initialized attention); our results generalize this to other time series FMs and use an even simpler supervised model (AR) to do so. Overall, these studies focus on specific cases, whereas our results identify a broader trend of FM underperformance and/or weak baselines across multiple fields.
Specialized baselines:
@@
Both automated supervised pipelines that we develop are heavily influenced by successful in-domain model development. In particular, the NAS-based pipeline we use to achieve our results in genomics and satellite imaging is inspired by the success of the human-driven specification of kernel sizes and dilation rates in successful architectures like TCN (Lea et al., 2016) and ConvNeXt (Liu et al., 2022). At the same time, for time series our approach is based upon a well-tuned GPU implementation of perhaps the most basic forecasting model, AR.
AutoML for specialized domains:
@@
While often evaluated on domains such as vision, automated techniques are also important in specialized domains. An important example is Auto-ARIMA (Hyndman & Khandakar, 2008) for time series, although it underperforms on the tasks we consider (Challu et al., 2022). However, to avoid requiring significant expertise in any one domain, we also make use of AutoML methods developed specifically for diverse tasks (Roberts et al., 2021b; Shen et al., 2023), in particular the NAS method DASH (Shen et al., 2022) that can discover good kernel sizes and dilation rates for a CNN backbone faster than it can be trained from scratch.
@@ 3 Methodology
While it is standard to automate the last two steps of the procedure, model development is typically done by hand and so is difficult to do for fifty tasks across three domains. As a result, we settle for approximating the traditional supervised workflow by simulating the model development component via neural architecture search. To ensure fair comparison and reduce computation, we use low-fidelity NAS methods that return an architecture in less time than it takes to train it. Our results can thus be viewed as lower bounds on the performance of supervised learning, as model development might be significantly improved using less-heuristic and/or human-driven architecture design.
In the remainder of this section we detail how we handle the different steps of the supervised pipeline. Note that our NAS-dependent workflow (DASHA)—which we cover in the first part of this section—yields our main results for genomics and satellite imaging but not for time series. There we find its performance to be less competitive and instead focus on an even simpler approach based on linear auto-regression, whose model development and tuning we describe in the second subsection.
@@ 3.1 DASHA: Simulating the supervised workflow using NAS
As described in Algorithm 1, we augment DASH in two ways: (1) trying more than one CNN backbone (e.g. both wide ResNet and UNet (Ronneberger et al., 2015)) and (2) using the popular hyperparameter optimization (HPO) method ASHA (Li et al., 2020) to configure training settings for the discovered model. Following the NAS and tuning stages, we train the discovered architecture with the selected configuration on the target data. Further details, including the resources given to the three steps of the pipeline and the exact search spaces used by DASH and ASHA, are provided in Appendix B.1. Note that, while our focus is on data-efficient baselines, we do ensure that the entire workflow is never substantially more computationally expensive than fine-tuning an FM.
@@ 4.1 Genomics
We begin our investigation in the genomics domain, which has witnessed the development of numerous FMs, including the early Enformer (Avsec et al., 2021), the DNABERT series (Ji et al., 2021; Zhou et al., 2023b), the HyenaDNA family (Nguyen et al., 2024), GENA-LM [^3] We compare to GENA-LM in Appendix C, as it reports metrics that differ from those of the NT benchmark. (Fishman et al., 2024), the recent Caduceus family (Schiff et al., 2024), and the NT family (Dalla-Torre et al., 2023); The latter includes models with up to 2.5B parameters. To evaluate them, we consider the Nucleotide Transformer (NT) benchmark of Dalla-Torre et al. (2023), which contains eighteen tasks in three main categories: regulatory elements, RNA production, and histone modification. We use this benchmark because of its diversity and because it has been evaluated on by all of the aforementioned FMs, allowing us to include eight of them in the comparison. Our numbers for these models are taken from Dalla-Torre et al. (2023, Supplementary Table 6); Following Dalla-Torre et al. (2023, Supplementary Table 5), We use F1 score and accuracy to evaluate a subset of regulatory elements and RNA production tasks, and we use Matthew’s Correlation Coefficient (MCC) as the main metric for evaluation on the remaining datasets.
Baselines:
CNNs have long been used for genomics tasks (Avsec et al., 2020; Cheng et al., 2024) and so are natural supervised baselines; in particular we include 1D variants of Wide ResNet (WRN) and UNet, which we find perform better than some domain-specific CNNs. We use these same two backbones as the candidate CNNs tuned and selected from by our DASHA workflow.
Model Model Pretraining Average Average Mean Median
@@
Foundation NT-Multispecies (2.5B) 2.5B 174B 0.697 4.08 40.51 45.52
Models DNABERT-2 117M 32.5B 0.680 6.88 38.65 43.59
@@
Supervised Wide ResNet 2.0M 0 0.694 6.83 37.16 43.08
Models UNet 4.5M 0 0.68 7.78 38.67 42.69
@@
Table 1: Aggregate genomics performance, showing that our supervised workflow (DASHA) attains state-of-the-art on the NT benchmark, outperforming all FMs according to all measures while using no pretraining data and oftentimes many fewer parameters. For Mean / Median %Imp., we report % improvement over the Raw Probe baseline from Dalla-Torre et al. (2023), and for DASHA the model size refers to the largest configuration across tasks..
Results:
Our genomics results are displayed in Table 1, which shows that our supervised workflow (DASHA) consistently outperforms all FMs across all aggregate metrics. As discussed in Appendix C, our strong performance is driven in large part by outstanding performance on the histone modification tasks (c.f. Table 9). The more detailed results also highlight the importance of considering diverse baselines, with Wide ResNet usually being the selected architecture but UNet performing significantly better for promoter and splice site classification tasks. Overall, DASHA sets a new state-of-the-art on the NT benchmark and certainly demonstrates that supervised methods remain quite competitive in genomics, despite the availability of massive pretraining datasets.
@@ 4.2 Satellite imaging
While they do not get as large as those in genomics, numerous BERT-scale FMs have also been introduced for satellite imaging, including SeCo (Manas et al., 2021), SatMAE (Cong et al., 2022), CROMA (Fuller et al., 2023), GFM (Mendieta et al., 2023), Scale-MAE (Reed et al., 2023), Satlas (Bastani et al., 2023), Prithvi (Jakubik et al., 2023), DOFA (Xiong et al., 2024), Clay (Clay Foundation, 2023), and SkySense (Guo et al., 2024). Because our evaluation includes GeoBench (Lacoste et al., 2024), a recently introduced satellite benchmark that has not been considered by many of these FMs, we obtain all results using our own fine-tuning; therefore we only consider a subset of top-performing open-source models. In all cases we use the fine-tuning workflow suggested by the authors of each FM plus some automated hyperparameter tuning; note that even with the extra tuning our reproductions on previous benchmarks systematically underperformed results reported in the original works. We select tasks mainly from GeoBench’s five classification tasks and then add four additional tasks—BigEarthNet (Sumbul et al., 2019), EuroSAT (Helber et al., 2019), Canadian Cropland (Jacques et al., 2023), and fMoW-Sentinel (Cong et al., 2022)—that are commonly used to evaluate other FMs. [^4] In Appendix C we report results when excluding tasks where missing channels may affect performance. As we focus on classification, we report top-1 accuracy or mAP as appropriate.
@@
SatMAE-Base 85.6M 700K 76.99 9.89 5.27 3.59
SatMAE-Large 303M 700K 77.75 7.56 6.52 4.62
GFM 86.8M 1.3M 77.18 9.00 5.77 4.08
SwinT-Base 86.8M 14M 76.69 8.72 4.86 1.43
Foundation CROMA-Base 90.6M 2M 77.39 7.22 5.85 4.22
Models CROMA-Large 312M 2M 78.03 5.44 6.90 6.09
DOFA-Base  [^1] M 8M 78.44 4.50 7.71 6.23
DOFA-Large 337M 8M 78.80 3.50 8.40 6.88
Satlas 90M 45M 77.47 6.33 5.87 3.86
ScaleMAE 323M 363.6K 77.10 8.21 5.42 4.40
Clay 92M 70M 77.33 7.50 5.99 3.51
ResNet50 23.5M 0 73.76 13.22 0.30 00.07
Supervised Wide ResNet 17.2M 0 73.97 12.89 0.00 0.00
Models UNet 17.3M 0 75.73 9.94 3.01 1.07
DASHA (our workflow) 32.4M 0 77.85 6.06 6.67 5.16
Table 2: Performance on satellite imaging, demonstrating that a supervised workflow (DASHA) can get within 1% average accuracy of state-of-the-art specialized FMs while using no pretraining data and having 2-10$\times$ fewer parameters. For Mean / Median %Imp. we report % improvement over a vanilla Wide ResNet, and for DASHA the model size refers to the largest configuration across tasks.
Baselines:
@@
Results:
Table 2 shows that our supervised workflow attains competitive performance across all aggregate metrics and is only slightly outperformed by CROMA-large and the DOFA family. Unlike in genomics, the FMs here consistently outperform CNN backbones, likely because the associated papers compare to them as baselines. However, the frequently superior performance of DASHA suggests that domain-aware model development would yield good supervised models in this field. Another contrast with genomics is that the larger FM versions consistently attain superior performance here, suggesting they are making at least somewhat effective use of the pretraining data. Nevertheless, that this improvement can also be attained by DASHA, which uses no pretraining and produces a model that is 3-10$\times$ smaller, suggests that there remains significant room for improvement.